For a projectile thrown into space with a speed v, the horizontal range is ((3)^(1/2)v^2)/2g. The vertical range is v^2/8g. The angle which the projectile makes with horizontal initially is
(a)15° (b)30° (c)45° (d)60°

Grade:10

2 Answers

Arun
5 years ago
Dear Adyasha
 
horizontal range = u^2 sin2@ / g
 
hence
 
sin2@ = sqrt(3)/ 2 
 
2@ = 60degree
 
@ = 30 degree
 
You can also check by height
Khimraj
5 years ago
max heightis given by
H = v2sin2\Theta/2g = v2/8g
so sin\Theta = ½
then \Theta = 30o
Hope it clears.............................